OpenStudy (anonymous):

What is the y-intercept of the function y = 10^x?

OpenStudy (btaylor):

it is in slope intercept form (y = mx + b), where (0, b) is the y-intercept. Here, b=0, so the y-intercept = (0,0)

OpenStudy (anonymous):

If x=0, then y = 10^0=1 (0,1) is the y-intercept.